OpenStudy (anonymous):

Solve for x: 3x − 5 = 2x + 6 1 −1 11 −11

OpenStudy (diyadiya):

Subtract 2x from both sides what do you get now?

OpenStudy (anonymous):

3x − 5 = 2x + 6 3x = 2x + 11 x = 11

OpenStudy (anonymous):

it was 11